Allen and Joy spent the same amount of money which was less than Php 150.00 for the toys they bought. Allen bought toys worth Php 15.00 each while Joy bought toys costing Php 20.00 each. What could be the largest amount that each spent for the toys?
Answer by Theo(13342)

basically, you're looking for the largest common multiple less than 150.

take 150 and divide it by 15 to get 10.
the largest number of toys allen could buy is 9.

take 150 and divide it by 20 to get 7.5.
the largest number of toys joy could buy is 7.

9 * 15 = 135.
7 * 20 = 140.

not the same.

8 * 15 = 120
6 * 20 = 120

they're the same !!!!!

the largest amount that each spent for the toys is 120.

the problem doesn't really state whether each spent 150 or they both spent 150 together.

if they both spent 150 together, then the problem is different.

look at multiples of 15 for 15 for allen and 20 for joy.

allen:
15,30,45,60,75,90,110...
joy:
20,40,60,80,100,120....

the largest amount that they both spent was 120.
that would be 60 apiece.

in the first rendition of this problem, it was assumed that they each spent less than 150.

in the second rendition of this problem, it was assumed that they both spent a total that was less than 150.

so your answer is that each spent either 120 or 60 depending on how you interpret th eproblem.